# Moving Mesh with Streamline Upwind Petrov-Galerkin (MM-SUPG) Method for Time-dependent Convection-Dominated Convection-Diffusion Problems

### v3

@@ 4.3 Temporal estimates
$M_{ij}$ $=\sum_{K\in\Omega_{i}\cap\Omega_{j}}\left(\frac{|K|}{(d+1)(d+2)}+\tau_{K}\int_{K}\varphi_{j}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}\right)$ (107)
The second integral term can be treated similar as (95). Now, we obtain the upper bound for $M_{ij}$,
$M_{ij}\leq\sum_{K\in\Omega_{i}\cap\Omega_{j}}\frac{|K|}{h_{i}^{K}[(d+1)(d+2)]^{3/2}}\left(h_{i}^{K}\sqrt{(d+1)(d+2)}+\tau_{K}\sqrt{2}(d+1)(d+2)\|{\boldsymbol{b}}\|_{\infty,K}\right).$ (108)
Recall we have the following bound on $A_{ij}$
$\begin{split}A_{ij}&\leq\sum_{K\in\Omega_{i}\cap\Omega_{j}}\frac{|K|\|{\boldsymbol{b}}\|_{\infty,K}}{h_{j}^{K}}\left(\frac{2}{\sqrt{(d+1)(d+2)}}+\frac{\tau_{K}\|{\boldsymbol{b}}\|_{\infty,K}}{h_{i}^{K}}\right)-\frac{\rho}{h_{i}^{K}h_{j}^{K}}\left(\tau_{K}\cos(\check{\alpha}_{ij})+|K|\cos(\tilde{\alpha}_{ij})\right)\\ &=\sum_{K\in\Omega_{i}\cap\Omega_{j}}\frac{1}{h_{i}^{K}h_{j}^{K}\sqrt{(d+1)(d+2)}}\left(2h_{i}^{K}h_{j}^{K}|K|\|{\boldsymbol{b}}\|_{\infty,K}+h_{j}^{K}|K|\tau_{K}\sqrt{(d+1)(d+2)}\|{\boldsymbol{b}}\|_{\infty,K}^{2}\right.\\ &\hskip 198.7425pt\left.-\rho\sqrt{(d+1)(d+2)}\left(\tau_{K}\cos(\check{\alpha}_{ij})+|K|\cos(\tilde{\alpha}_{ij})\right)\right)\end{split}$ (109)
Thus, $\tilde{\boldsymbol{A}}$ in (33) is a $Z$-matrix using the assumption on $\Delta t$.
By (33), we can write
$\tilde{\boldsymbol{A}}=\begin{bmatrix}M_{11}+\theta\Delta tA_{11}&M_{12}+\theta\Delta tA_{12}\\ 0&I_{N_{v}-N_{vi}}\end{bmatrix}.$
Since $\tilde{\boldsymbol{A}}$ is a $Z$-matrix, the block $M_{11}+\theta\Delta tA_{11}$ is also a $Z$-matrix. First, we show $M_{11}$ is positive-definite. Let ${\boldsymbol{v}}=(v_{1},v_{2},\dots,v_{N_{vi}})$ be a nonzero vector in $\mathbb{R}^{N_{vi}}$ and write $v^{h}=\sum_{j=1}^{N_{vi}}v_{j}\varphi_{j}\in U_{h}^{0}$. Using the assumption on the divergence of ${\boldsymbol{b}}$ being either nonnegative of $\tau_{K}$ satisfying (89), ${\boldsymbol{v}}^{T}M_{11}{\boldsymbol{v}}>0$ and ${\boldsymbol{v}}^{T}A_{11}{\boldsymbol{v}}>0$ (this is a similar calculation as shown in Lemma 4.4). Therefore, $M_{11}+\theta\Delta tA_{11}$ is an $M$-matrix. We have
$\tilde{\boldsymbol{A}}^{-1}=\begin{bmatrix}(M_{11}+\theta\Delta A_{11})^{-1}&-(M_{11}+\theta\Delta tA_{11})^{-1}(M_{12}+\theta\Delta tA_{12})\\ 0_{N_{v}}&I_{N_{v}-N_{vi}}\end{bmatrix},$ (110)
which implies $\tilde{\boldsymbol{A}}$ is an $M$-matrix since $\tilde{\boldsymbol{A}}^{-1}\geq 0$ (note that $M_{12}+\theta\Delta tA_{12}\leq 0$). ∎
@@
(104) holds along the diagonal entries if the mesh generated by $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$ satisfy ANAC and are of SPD-type, and $\Delta t$ satisfies the bound
$\Delta t\leq\frac{1}{(1-\theta)(d+1)(d+2)}\min_{K\in\mathcal{T}_{h}}\min_{i,j=1,\dots,d+1}\frac{(\tilde{h}_{i}^{K}\check{h}_{i}^{K})^{2}(\tau_{K}(d+1)(d+2)\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})+2)}{|K|(\check{h}_{i}^{K})^{2}+(\tilde{h}_{i}^{K}\check{h}_{i}^{K})^{2}(\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})+\tau_{K}({\boldsymbol{b}}\cdot\nabla\varphi_{i})^{2})+\tau_{K}(\tilde{h}_{i}^{K})^{2}}$ (111)
@@
From (88) and (81), along with (54), it’s clear that if (111) holds, then $M_{ii}-(1-\theta)A_{ii}\geq 0$. ∎
@@ 4.4 Metric tensor for DMP
Suppose the two meshes generated by $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$ satisfy ANAC, are of SPD-type, and satisfy the step size bounds (105) and (111). If ${\boldsymbol{b}}$ has nonpositive divergence or $\tau_{K}$ satisfies (89), then the SUPG scheme satisfies DMP.
Since two different metric tensors will be used on any given node, we follow [40] and taken into consideration for every time step, we use metric intersection in order to generate the anisotropic mesh.
@@
[…]
@@ 4.5 Numerical Results
In this section we present our numerical findings for using different metric tensors. We start with the metric tensors $\mathcal{M}_{DMP}=\theta\mathbb{D}^{-1}$ and $\mathcal{M}_{\mathbb{K}}=\theta^{\prime}\mathbb{K}^{-1}$ and define the metric tensor $\mathcal{M}_{int}=\mathcal{M}_{DMP}\cap\mathcal{M}_{\mathbb{K}}$. We use the implicit Euler method in time for all computations (that is, $\theta=1$) and compare the results with the classic Galerkin method with SUPG using fixed mesh (FM-SUPG).
@@
Consider the manufactured solution on the domain $\Omega_{T}=[0,1]^{2}\times[0.2,0.3]$ to (1) given by
$u(x,y,t)=(1+\exp(C(x+y-t))^{-1},\ \ C>0$ (116)
@@
The diffusion tensor has eigenvalues $\lambda_{1}=100$ and $\lambda_{2}=1$, which indicates that the diffusion in horizontal direction is 100 times faster than that in the vertical direction. For the given flow in (118), the tensor $\mathbb{K}$ defined in (50) is the same as $\mathbb{D}$. Therefore, the metric tensor $\mathcal{M}_{DMP}$ is the same as $\mathcal{M}_{\mathbb{K}}$ for this example.
Table 5 presents the minimum values of the numerical solutions obtained from different meshes. As can be seen, the results from $\mathcal{M}_{int}$ and $\mathcal{M}_{DMP}$ are nearly identical. Even the results from the fixed mesh show satisfaction of DMP for the bottom shelf. However, the graph in Figure 13(a) shows significant artificial oscillation near the top shelf when using the fixed mesh approach. On the other hand, as shown in Figure 13(b), when using $\mathcal{M}_{int}$ or $\mathcal{M}_{DMP}$, the artificial oscillation at the top shelf is essentially eliminated.
Table 5: Example 4.1. $\underset{\Omega_{T}}{\min}\ u$ with $\Delta t=10^{-2}$.
$N$ $\mathcal{M}_{int}$ $\mathcal{M}_{DMP}$ Fixed mesh
16200 -4.4374e-05 -4.4373e-05 -2.8664e-07
24200 -3.1365e-06 -3.1367e-06 -2.2583e-08
34322 -1.5736e-07 -1.5735e-07 -1.1880e-09
@@
Figure 13: Example 4.1 – side view of solutions for a fixed mesh and $\mathcal{M}_{int}$ with $N=34322$, $\varepsilon=10^{-6}$, and $\Delta t=10^{-2}$.
@@
$\mathbb{D}=\begin{pmatrix}1&1\\ 1&3\end{pmatrix}$ (119)
@@
To see the advantage of $\mathcal{M}_{DMP+adapt}$ mesh over $\mathcal{M}_{int+adapt}$ mesh, we compute the numerical solutions at different $N$ values with larger time step $\Delta t=10^{-3}$. The minimum values of the numerical solutions are summarized in Table 6. As can be seen, $\mathcal{M}_{int+adapt}$ provides the best results.
Table 6: Example 4.2. Results for $\mathcal{M}_{int+adapt}$, $\mathcal{M}_{DMP+adapt}$, and fixed mesh for varying $N$ with $\varepsilon=10^{-3}$ and $\Delta t=10^{-3}$.
$N$ Fixed Mesh $\mathcal{M}_{DMP+adapt}$ $\mathcal{M}_{int+adapt}$

### v4

@@ 4.3 Temporal estimates
$M_{ij}=\sum_{K\in\omega_{i}\cap\omega_{j}}\left(\frac{|K|}{(d+1)(d+2)}+\tau_{K}\int_{K}\varphi_{j}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}\right).$ (102)
Applying (49), we have
$M_{ij}\leq\sum_{K\in\omega_{i}\cap\omega_{j}}\left(\frac{|K|}{(d+1)(d+2)}+\tau_{K}\frac{|K|}{(d+1)h_{i}^{K}}\,\|b\|_{\infty,K}\right)=\sum_{K\in\omega_{i}\cap\omega_{j}}\mu_{ij}^{K},$ (103)
and from (93), we obtain
$A_{ij}\leq-\sum_{K\in\omega_{i}\cap\omega_{j}}\eta_{ij}^{K}.$ (104)
Thus,
$\tilde{\boldsymbol{A}}_{ij}=M_{ij}+\theta\Delta tA_{ij}\leq\sum_{K\in\omega_{i}\cap\omega_{j}}\mu_{ij}^{K}-\theta\Delta t\sum_{K\in\omega_{i}\cap\omega_{j}}\eta_{ij}^{K}\leq 0$ (105)
provided (101) holds. Hence all off-diagonal entries of $\tilde{\boldsymbol{A}}$ are nonpositive. Combined with the positivity of the diagonal entries of $\boldsymbol{M}$} and $\boldsymbol{A}$}, $\tilde{\boldsymbol{A}}$ is a $Z$-matrix.
Next, we show that $\tilde{\boldsymbol{A}}$ is an $M$-matrix. Reorder the system into interior and boundary nodes, by (33), we can write
$\tilde{\boldsymbol{A}}=\begin{bmatrix}\boldsymbol{B}&\boldsymbol{C}\\ 0&\boldsymbol{I}\end{bmatrix},\quad\boldsymbol{B}:=\boldsymbol{M}_{11}+\theta\Delta t\boldsymbol{A}_{11}.$
Since $\boldsymbol{C}\leq 0$, it suffices to show that $\boldsymbol{B}$ is an $M$-matrix.
Firstly, the block $\boldsymbol{B}$ is also a $Z$-matrix because $\tilde{\boldsymbol{A}}$ is a $Z$-matrix. Next, we show $\boldsymbol{B}$ is positive-definite.
Let ${\boldsymbol{v}}=(v_{1},v_{2},\dots,v_{N_{vi}})$ be a nonzero vector in $\mathbb{R}^{N_{vi}}$ and write $v_{h}=\sum_{j=1}^{N_{vi}}v_{j}\varphi_{j}\in U_{0}^{h}$. Then,
$\boldsymbol{v}^{T}\boldsymbol{B}\boldsymbol{v}=\boldsymbol{v}^{T}\boldsymbol{M}_{11}\boldsymbol{v}+\theta\Delta t\boldsymbol{v}^{T}\boldsymbol{A}_{11}\boldsymbol{v},$ (106)
and the interior mass bloack satisfies
$\boldsymbol{v}^{T}\boldsymbol{M}_{11}\boldsymbol{v}=\sum_{i,j=1}^{N_{vi}}v_{i}M_{ij}v_{j}.$ (107)
Using the definition of $M_{ij}$ in (26), this becomes
$\boldsymbol{v}^{T}\boldsymbol{M}_{11}\boldsymbol{v}=\sum_{K\in\mathcal{T}_{h}}\left(\int_{K}v_{h}^{2}\textrm{d}\boldsymbol{x}+\tau_{K}\int_{K}v_{h}(\boldsymbol{b}\cdot\nabla v_{h})\textrm{d}\boldsymbol{x}\right).$ (108)
Applying Cauchy-Schwarz and inverse inequality, we have
$\left|\int_{K}v_{h}(\boldsymbol{b}\cdot\nabla v_{h})\textrm{d}\boldsymbol{x}\right|\leq C_{inv}h_{K}^{-1}\|b\|_{\infty,K}\|v_{h}\|^{2}_{L^{2}(K)}.$ (109)
Therefore,
[…]
$=\sum_{K}\left(1-C\tau_{K}h_{K}^{-1}\|b\|_{\infty,K}\right)\|v_{h}\|^{2}_{L^{2}(K)}$
$>0,$
provided $\tau_{K}$ satisfies the condition (99).
Similarly, from (27) and (60), we get,
$\boldsymbol{v}^{T}\boldsymbol{A}_{11}\boldsymbol{v}$ $=\sum_{K\in\mathcal{T}_{h}}\left(\int_{K}v_{h}(\boldsymbol{b}\cdot\nabla v_{h})\textrm{d}\boldsymbol{x}+\tau_{K}\int_{K}(\boldsymbol{b}\cdot\nabla v_{h})^{2}\textrm{d}\boldsymbol{x}\right)$ (110)
$+\sum_{K\in\mathcal{T}_{h}}\varepsilon|K|(\nabla v_{h})^{T}(\mathbb{D}_{K}+\tau_{K}\mathbb{K}_{K})\nabla v_{h}.$
Applying Young’s inequality to the scaler quantities $v_{h}$ and $\boldsymbol{b}\cdot\nabla v_{h}$, we have
$v_{h}(b\cdot\nabla\phi_{i})+\tau_{K}(b\cdot\nabla\phi_{i})^{2}\geq-\frac{1}{4\tau_{K}}v_{h}^{2}.$ (111)
Integrate over $K$ and substitue into (110), we have
$\boldsymbol{v}^{T}\boldsymbol{A}_{11}\boldsymbol{v}$ $\geq\sum_{K\in\mathcal{T}_{h}}\left(-\frac{1}{4\tau_{K}}\|v_{h}\|^{2}_{L^{2}(K)}+\varepsilon|K|(\nabla v_{h})^{T}(\mathbb{D}_{K}+\tau_{K}\mathbb{K}_{K})\nabla v_{h}\right).$
Using SPD properties of $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$, we obtain
$(\nabla v_{h})^{T}(\mathbb{D}_{K}+\tau_{K}\mathbb{K}_{K})\nabla v_{h}\geq(\rho_{min}(\mathbb{D}_{K})+\tau_{K}\rho_{min}(\mathbb{K}_{K}))|\nabla v_{h}|^{2}.$ (112)
Since $v_{h}$ is linear on each element $K$, $\nabla v_{h}$ is constant on $K$, so
$|K||\nabla v_{h}|^{2}=\|\nabla v_{h}\|^{2}_{L^{2}(K)}.$ (113)
Therefore,
[…]
Applying inverse inequality
$\|v_{h}\|_{L^{2}(K)}\leq C_{inv}h_{K}\|\nabla v_{h}\|^{2}_{L^{2}(K)},$ (114)
we have
$\boldsymbol{v}^{T}\boldsymbol{A}_{11}\boldsymbol{v}\geq\sum_{K}\left(\varepsilon(\rho_{min}(\mathbb{D}_{K})+\tau_{K}\rho_{min}(\mathbb{K}_{K}))-C^{2}_{inv}\frac{h_{K}^{2}}{4\tau_{K}}\right)\|\nabla v_{h}\|^{2}_{L^{2}(K)}>0,$ (115)
provided the condition (100) holds. Thus, $\boldsymbol{v}^{T}\boldsymbol{B}\boldsymbol{v}>0$ for all $\boldsymbol{v}\neq\boldsymbol{0}$. So $\boldsymbol{B}$ is positive definite.
Since $\boldsymbol{B}$ is a $Z$-matrix that is positive definite, it follows that $\boldsymbol{B}$ is a nonsingular $M$-matrix. Hence $\boldsymbol{B}^{-1}\geq 0$. Then,
$\tilde{\boldsymbol{A}}^{-1}=\begin{bmatrix}\boldsymbol{B}^{-1}&-\boldsymbol{B}^{-1}\boldsymbol{C}\\ 0&\boldsymbol{I}\end{bmatrix}\geq 0,$
since $\boldsymbol{C}\leq 0$. Therefore, $\tilde{\boldsymbol{A}}$ is an $M$-matrix.
Under the same assumptions in Lemma 4.4, we have
$M_{ii}\geq\sum_{K\in\omega_{i}}\left(\frac{2|K|}{(d+1)(d+2)}-\tau_{K}\frac{|K|}{(d+1)h_{i}^{K}}\|b\|_{\infty,K}\right),$
and
$A_{ii}$ $=\sum_{K\in\omega_{i}}\int_{K}\left(\varphi_{i}{\boldsymbol{b}}\cdot\nabla\varphi_{i}+\tau_{K}({\boldsymbol{b}\cdot\nabla\varphi_{i}})^{2}\right)\textrm{d}{\boldsymbol{x}}+|K|(\nabla\varphi_{i})^{T}\varepsilon(\tau_{K}\mathbb{K}_{K}+\mathbb{D}_{K})\nabla\varphi_{i}$
$\leq\sum_{K\in\omega_{i}}|K|\left(\frac{\|b\|_{\infty,K}}{(d+1)h_{i}^{K}}+\frac{\|b\|^{2}_{\infty,K}}{(h_{i}^{K})^{2}}+\varepsilon\frac{\rho_{min}(\mathbb{D}_{K})+\tau_{K}\rho_{min}(\mathbb{K}_{K})}{(h_{i}^{K})^{2}}\right).$ (116)
Then we have the following lemma.
@@
Assume
$M_{ii}^{low}:=\sum_{K\in\omega_{i}}\left(\frac{2|K|}{(d+1)(d+2)}-\tau_{K}\frac{|K|}{(d+1)h_{i}^{K}}\|b\|_{\infty,K}\right)>0,$ (117)
and let
$A_{ii}^{up}:=\sum_{K\in\omega_{i}}|K|\left(\frac{\|b\|_{\infty,K}}{(d+1)h_{i}^{K}}+\frac{\|b\|^{2}_{\infty,K}}{(h_{i}^{K})^{2}}+\varepsilon\frac{\rho_{min}(\mathbb{D}_{K})+\tau_{K}\rho_{min}(\mathbb{K}_{K})}{(h_{i}^{K})^{2}}\right).$ (118)
If
$\Delta t\leq\frac{1}{1-\theta}\min_{i}\frac{M_{ii}^{low}}{A_{ii}^{up}},$ (119)
then
$M_{ii}-\Delta t(1-\theta)A_{ii}\geq 0\quad\text{ for all }i.$ (120)
@@
Since $M_{ii}\geq M_{ii}^{low}$ and $A_{ii}\leq A_{ii}^{up}$, we have
$M_{ii}-\Delta t(1-\theta)A_{ii}\geq M_{ii}^{low}-\Delta t(1-\theta)A_{ii}^{up}\geq 0$ (121)
whenever (119) holds. ∎
@@ 4.4 Metric tensor for DMP
Consider the fully discrete MM-SUPG scheme (26)-(28). Assume that, for each element $K\in\mathcal{T}_{h}$, the following conditions hold:
  1. (i)
    The elementwise tensors $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$ are symmetric positive definite.
  2. (ii)
    The meshes associated with $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$ satisfy the anisotropic nonobtuse angle condition (ANAC).
  3. (iii)
    The stabilization parameter $\tau_{K}$ satisfies (86), (87), and (90).
  4. (vi)
    The time step $\Delta t$ satisfies (101) and (119).
Then the coefficient matrices satisfy:
  • (a)
    $\boldsymbol{M}-\Delta t(1-\theta)\boldsymbol{A}$ has nonnegative entries,
  • (b)
    $\widetilde{\boldsymbol{A}}=\boldsymbol{M}+\theta\Delta t\boldsymbol{A}$ is an $M$-matrix.
Consequently, the fully discrete MM-SUPG scheme satisfies the discrete maximum principle (DMP), i.e., the numerical solution preserves the bounds of the initial and boundary data.
Proof.
The results follows immediately from the previous lemmas. ∎
Corollary 4.10.1.
Under the assumptions of Theorem 4.10, the discrete maximum principle (DMP) holds if the following practical conditions are satisfied:
  1. (i)
    Mesh condition: The mesh is sufficiently aligned with the anisotropy of the problem in the sense that it satisfies the anisotropic nonobtuse angle condition (ANAC) with respect to both $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$.
  2. (ii)
    Stabilization condition: The stabilization parameter $\tau_{K}$ is chosen such that
    $\tau_{K}\|b\|_{\infty,K}\lesssim h_{K},$
    where $h_{K}$ denotes a representative element size.
  3. (iii)
    Balance condition: The combined diffusion and stabilization effects dominate the convection contribution, i.e.,
    $\rho_{\min}(\varepsilon\mathbb{D}_{K})+\tau_{K}\rho_{\min}(\varepsilon\mathbb{K}_{K})\gtrsim h_{K}\|b\|_{\infty,K}+\tau_{K}\|b\|_{\infty,K}^{2}.$
  4. (iv)
    Time step condition: The time step $\Delta t$ is chosen within a suitable range such that
    $\Delta t\gtrsim\frac{1}{\theta}\frac{M_{ij}}{|A_{ij}|}\quad\text{and}\quad\Delta t\lesssim\frac{1}{1-\theta}\frac{M_{ii}}{A_{ii}},$
    ensuring that $\widetilde{A}=M+\theta\Delta tA$ is an $M$-matrix and that $M-\Delta t(1-\theta)A$ remains nonnegative.
Under these conditions, the fully discrete MM-SUPG scheme satisfies the discrete maximum principle.
Remark 1.
The corollary highlights that the DMP is achieved through a balance between mesh alignment, stabilization, and time stepping. In particular, the interaction between the mesh metric and the SUPG stabilization plays a key role in controlling oscillations and preserving monotonicity.
Remark 2.
Since we use implicit Euler method in time for all computations (that is, $\theta=1$), the lower bound (101) for time step $\Delta t$ is practically removed.
Since two different metric tensors will be used on any given node, we follow [42] and take into consideration for every time step, we use metric intersection in order to generate the anisotropic mesh.
@@
$\tilde{K}Kˇ\check{K}$$\mathbb{D}_{K}^{-1/2}{\boldsymbol{x}}$$\mathbb{K}_{K}^{-1/2}{\boldsymbol{x}}$
@@ 4.5 Numerical Results
In this section we present our numerical findings for using different metric tensors. We start with the metric tensors $\mathcal{M}_{DMP}=c_{1}\mathbb{D}^{-1}$ and $\mathcal{M}_{\mathbb{K}}=c_{2}\mathbb{K}^{-1}$, where $c_{1}$ and $c_{2}$ are arbitray constants. Define the metric tensor
$\mathcal{M}_{int}=\mathcal{M}_{DMP}\cap\mathcal{M}_{\mathbb{K}}.$ (126)
We use the implicit Euler method in time for all computations (that is, $\theta=1$) and compare the results with the classic Galerkin method with SUPG using fixed mesh (FM-SUPG).
@@
Consider the manufactured solution on the domain $\Omega_{T}=[0,1]^{2}\times[0,0.3]$ to (1) given by
$u(x,y,t)=(1+\exp(C(x+y-t))^{-1},\ \ C>0,$ (127)
@@
The diffusion tensor has eigenvalues $\lambda_{1}=100$ and $\lambda_{2}=1$, which indicates that the diffusion in horizontal direction is 100 times faster than that in the vertical direction. For the given flow in (129), the tensor $\mathbb{K}$ defined in (58) is the same as $\mathbb{D}$. Therefore, the metric tensor $\mathcal{M}_{DMP}$ is the same as $\mathcal{M}_{\mathbb{K}}$ for this example. Hence, for this example, $\mathcal{M}_{int}$ is also the same as $\mathcal{M}_{DMP}$.
Figure 12 shows the side view of the numerical solutions using different metric tensors. As can be seen, the graph in Figure 12(a) shows significant artificial oscillation near the top shelf when using the fixed mesh approach. On the other hand, as shown in Figure 12(b), when using $\mathcal{M}_{int}$ or $\mathcal{M}_{DMP}$, the artificial oscillation at the top shelf is essentially eliminated.
@@
Figure 12: Example 4.1 – side view of solutions (scaled-up) for a fixed mesh and $\mathcal{M}_{int}$ with $N=34322$, $\varepsilon=10^{-6}$, and $\Delta t=10^{-2}$.
@@
$\mathbb{D}=\begin{pmatrix}1&1\\ 1&3\end{pmatrix}$, (130)
@@
To see the advantage of $\mathcal{M}_{int+adapt}$ mesh over $\mathcal{M}_{DMP+adapt}$ mesh, we compute the numerical solutions at different $N$ values with larger time step $\Delta t=10^{-3}$. The minimum values of the numerical solutions are summarized in Table 5. As can be seen, except on the course mesh, $\mathcal{M}_{int+adapt}$ provides the best results.
Table 5: Example 4.2. Results for $\mathcal{M}_{int+adapt}$, $\mathcal{M}_{DMP+adapt}$, and fixed mesh for varying NN with $\varepsilon=10^{-3}$ and $\Delta t=10^{-3}$.
NN Fixed Mesh $\mathcal{M}_{DMP+adapt}$ $\mathcal{M}_{int+adapt}$